1. A reaction was shown to follow second-order kinetics. How much time is required for [A] to change from 0.500 M to 0.160 M? (k = 0.456 M⁻¹ s⁻¹)

2. A substance decomposes with a rate constant of 9.05 × 10⁻⁴ s⁻¹. How long does it take for 16.0% of the substance to decompose?

3. How long will it take for the concentration of A to decrease from 0.500 M to 0.100 M in the first-order reaction A → B? (k = 0.800 s⁻¹)
Answer #1

1)

we have:

[A]o = 0.5 M

[A] = 0.16 M

k = 0.456 M-1.s-1

use integrated rate law for 2nd order reaction

1/[A] = 1/[A]o + k*t

1/(0.16) = 1/(0.5) + 0.456*t

6.25 = 2 +0.456*t

0.456*t = 4.25

t = 9.32 s

Answer: 9.32 s

Only 1 question at a time please